# Libo Qin  Qiguang Chen  Xiachong Feng  Yang Wu  Yongheng Zhang Yinghui Li  Min Li  Wanxiang Che  Philip S. Yu

### v1

@@ 5.5 Hallucination in LLMs for NLP
During solving the NLP tasks, LLMs inevitably suffer from the hallucinations where LLMs produce outputs that deviate from world knowledge Muhlgay et al. (2023); Min et al. (2023), user request Adlakha et al. (2023), or self-generated context Liu et al. (2022). This deviation harms the reliability of LLMs in practical scenarios.
@@ 5.6 Safety in LLMs for NLP
Applying large models to downstream NLP tasks also raises inevitable safety concerns, including copyright issues Chang et al. (2023), hate toxicity Hartvigsen et al. (2022), social bias Wan et al. (2023a); Dhamala et al. (2021) and psychological safety Huang et al. (2023b). Inspired by this, a series of works focus on the research on the safety of LLMs for diverse NLP tasks Ganguli et al. (2022); Sun et al. (2023a).
The main challenges to safety in LLMs are: (1) Safety Benchmark Construction: Currently, there are few security-related benchmarks for LLM on various NLP tasks. Establishing effective safety benchmarks is a critical objective in this area. (2) Multilingual Safety Risks: LLM suffers more safety risks across languages and cultures. Identifying and mitigating these risks in a multilingual context is a significant challenge.
@@ 6 Conclusion
In this work, we make the first attempt to offer a systemic overview of LLMs in NLP, introducing a unified taxonomy about parameter-frozen applications and parameter-tuning applications. Besides, we highlight new research frontiers and challenges, hoping to facilitate future research. Additionally, we maintain a publicly available resource website to track the latest developments in the literature. We hope this work can provide valuable insights and resources to build effective LLMs in NLP.

### v2

@@ 5.5 Hallucination in LLMs for NLP
During solving the NLP tasks, LLMs inevitably suffer from the hallucinations where LLMs produce outputs that deviate from world knowledge [292, 293], user request [294], or self-generated context [295]. This deviation harms the reliability of LLMs in practical scenarios.
@@ 5.6 Safety in LLMs for NLP
Applying large models to downstream NLP tasks also raises inevitable safety concerns, including copyright issues [296], hate toxicity [297], social bias [298, 299] and psychological safety [300]. Inspired by this, a growing body of research has emerged, focusing on ensuring the safety of LLMs for various NLP tasks [301, 302, 303, 304].
The main challenges to safety in LLMs are: (1) Safety Benchmark Construction: Currently, there are few security-related benchmarks for LLM on various NLP tasks. Establishing effective safety benchmarks is a critical objective in this area. (2) Multilingual Safety Risks: LLM suffers more safety risks across languages and cultures [305]. Identifying and mitigating these risks in a multilingual context is a significant challenge.
5.7 Long Chain-of-Thought in LLMs for NLP
Long Chain-of-Thought (Long-CoT) extends standard CoT prompting by allowing models to reason more deeply, explore multiple solution paths, and reflect on intermediate outcomes instead of following a single linear chain of thought [28, 306, 307]. By organizing reasoning into hierarchical levels or segmented sub-chains, Long-CoT equips large language models to address complex NLP challenges and compositional reasoning tasks beyond the reach of conventional CoT [308, 309, 310, 16, 311, 288, 312]. Recent innovations integrate reflective mechanisms [282, 313], inference-time scaling techniques [116, 314, 315], and reinforcement-learning enhancements [7, 316, 317, 318, 319].
Key challenges in this direction include: (1) Adaptive Reasoning Length Control: Selecting the appropriate depth and breadth for each sub‑chain is challenging [320, 321, 322]. If a sub‑chain is too shallow, the model may overlook critical intermediate abstractions; if it is too deep, it risks propagating errors or exceeding token limits [19, 323]. (2) Interactive Reasoning: Enabling a dynamic, iterative problem‑solving process, where models pose clarifying questions [324], integrate external feedback [281, 325], and refine intermediate steps [326], remains insufficiently explored [327, 28]. Such interactive chains could substantially improve performance and accuracy in tasks requiring real‑time adaptation [328, 329].
@@ 6 Conclusion
In this work, we make the first attempt to offer a systemic overview of LLMs in NLP, introducing a unified taxonomy of parameter-frozen paradigm and parameter-tuning paradigm. Besides, we highlight new research frontiers and challenges, hoping to facilitate future research. Additionally, we maintain a publicly available resource website to track the latest developments in the literature. We hope this work can provide valuable insights and resources to build effective LLMs in NLP.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (NSFC) via grant 62306342, 62236004, 62206078 and 62476073. This work was supported by the Scientific Research Fund of Hunan Provincial Education Department (24B0001). This work was sponsored by the Excellent Young Scientists Fund in Hunan Province (2024JJ4070), the Science and Technology Innovation Program of Hunan Province under Grant 2024RC3024 and CCF-Zhipu Large Model Innovation Fund (NO.CCF-Zhipu202406). This work was carried out in part using computing resources at the High Performance Computing Center of Central South University.